# HBAR entropy in causal diamond geometry: A near-horizon perspective

### v1

@@ Abstract
In this article we extend the Horizon Brightened Acceleration Radiation (HBAR) framework to the causal diamond (CD) spacetime. We study a cloud of two-level atoms, injected at random times in the asymptotic past of the CD, freely falling toward its causal horizon and emitting scalar radiation via a weak dipole coupling to a quantum field. In the near-horizon region, an emergent conformal symmetry—captured by conformal quantum mechanics (CQM)—governs the field dynamics and allows analytic control of the emission process. We find that the radiation spectrum is thermal, with temperature $T_{D}=1/(\pi\alpha)$, and that the associated von Neumann entropy flux reproduces the entropy production of the radiation field. These results demonstrate that the causal horizons of the CD spacetime effectively act as a topological thermal reservoir, with thermal properties arising entirely from the global causal structure rather than from underlying microscopic degrees of freedom, highlighting that the validity of the HBAR framework is fundamentally tied to the existence of causal horizons, independent of the presence of a black hole.
@@ I Introduction
In the context of Horizon Brightened Acceleration Radiation (HBAR) entropy, as proposed by Marlan Scully et al. in Ref. [1], it has been argued that the nature of the radiation emitted by a cloud of atoms—randomly injected from the asymptotic past and following free-falling (geodesic) trajectories towards the event horizon of a black hole—is a consequence of the relative acceleration between the field modes, defined on the black hole geometry with a well-defined Boulware vacuum state [2], and the locally inertial motion of the atoms. The origin of this radiation lies in the weak dipole coupling between the atoms, modeled as two-level systems, and the quantum field [3]. A characteristic process of this interaction is the excitation of an atom accompanied by the emission of a scalar quantum, which can either reach future timelike infinity—thereby contributing to the observable radiation spectrum—or be reabsorbed by another atom within the cloud before being emitted again.
@@
Consequently, the HBAR framework finds a natural extension in the near-horizon geometry, providing a powerful tool for analyzing more complex spacetime geometries, especially in situations where an exact geodesic analysis is intractable. The conformal symmetry that emerges in this regime allows for analytic control over the scalar quanta emission and absorption spectra associated with acceleration radiation. This extension has proven fruitful in subsequent analyses of HBAR entropy for generalized geometries, including Schwarzschild black holes [9], Kerr black holes [11, 12, 13], and even non-asymptotically flat spacetimes such as the BTZ geometry [14].
@@ II.1 Coordinate transformation and metric tensor
The spacetime experienced by such a finite-lifetime observer defines the so-called Causal Diamond (CD) spacetime. It can be geometrically characterized from the subset of Minkowski spacetime points restricted to the region $D_{R}$, denoted as $x^{\mu}=(t_{d_{R}},x_{d_{R}})$. These coordinates are obtained by applying the conformal transformation introduced in Ref. [15] to the Minkowski coordinates restricted to the right Rindler wedge $R$ (see Fig. 1(b)), which are given by $\tilde{x}^{\mu}=(t_{r},x_{r})$. Accordingly, the conformal map is defined as
@@
where $\tilde{\alpha}=2\alpha/\lambda$. Then, by applying Eqs. (1)–(5), we obtain the coordinate transformation between the Minkowski points restricted to the region $D_{R}$ and the CD spacetime coordinates $(\eta,\rho)$:
@@
The line element of the causal diamond (CD) spacetime takes the form
@@ II.2.1 Causal diamond horizons
We begin by analyzing the expansion of $(t_{d_{r}},x_{d_{r}})$ around the regime $\rho\ll 1$, which is given by:
@@
In addition, since $\sinh(2\eta/\alpha)\approx\pm\cosh(2\eta/\alpha)$ for $\eta\gg 1$ and $\eta\gg 1$, respectively, it follows from Eqs. (10) and (11) that, in the simultaneous limits $\rho\to 0$ and $\eta\to\pm\infty$, one directly obtains the following causal horizons (see Fig. 3(a)):
@@
On the other hand, we now turn to the expansion of $(t_{d_{r}},x_{d_{r}})$ in the regime $\rho\gg 1$, from which we obtain:
@@
As in the previous case, applying the limits $\rho\to\infty$ and $\eta\to\pm\infty$ to Eqs. (14) and (15) yields the following causal horizons (see Fig. 3(b)):
@@ II.2.2 Causal diamond corners
The asymptotic regions of the CD spacetime corresponding to the four corners of $D_{R}$ can be identified by analyzing the coordinate transformations between Minkowski and the CD spacetime in the appropriate limits:
@@ II.3 Effective surface gravity of the CD spacetime horizons
Likewise, noting that the boundary of the CD spacetime corresponds, in Minkowski spacetime, to the boundary of the region $D_{R}$, given by $\partial D_{R}:=\quantity{(t,x):\lvert t\rvert+\lvert x\rvert=\alpha}$, we observe that it precisely corresponds to the Killing horizon of the vector field $\chi^{(\eta)}$:
@@
Thus, the effective surface gravity associated with the conformal Killing vector $\chi^{(\eta)}$ is defined by evaluating the following identity on the boundary of $D_{R}$:
@@
where $\hat{\epsilon}=\pm 1$ depending on whether the normal vector to a given segment of the boundary of $D_{R}$ is future-directed $(+)$ or past-directed ()(-) (see Fig. 5(b)).
Finally, by evaluating Eq. (24) explicitly in Minkowski coordinates—using the expression for $\chi^{(\eta)}$ given in Eq. (22) together with the relation $\chi^{(\eta)}\cdot\chi^{(\eta)}=\eta_{ab}\chi^{a}_{(\eta)}\chi^{b}_{(\eta)}$—we obtain the following value for the effective surface gravity and its corresponding Hawking-like temperature:
@@
Thus, the obtained temperature exactly matches the temperature $T_{D}$ associated with the thermal description of the Minkowski vacuum as perceived by an observer with finite lifetime $2\alpha$, restricted to the causal diamond $D_{R}$. This result is consistent with previous findings based on the thermal time hypothesis, the open quantum system approach, the thermofield double construction, and related frameworks [18, 19, 20]. In what follows, we adopt this identification of $T_{D}$ in terms of the surface gravity $\kappa$, namely $T_{D}=\kappa/(2\pi)=(\pi\alpha)^{-1}$.
@@ II.4.1 Geodesic trajectories
From the preceding geometric analysis, we conclude that, when viewed from Minkowski spacetime, the causal diamond $D_{R}$ exhibits a structure analogous to that of a Penrose–Carter diagram. This correspondence follows from the fact that the CD spacetime metric, given in Eq. (8), is conformally related to the Minkowski metric through the coordinate transformation $\rho=e^{2\xi/\alpha}$.
@@
Figure 6: (a) Timelike geodesics. (b) A timelike geodesic in the extended representation of $i^{-}$ and $i^{+}$. (b) Null geodesics.
@@ II.4.2 Geodesics in the near-horizon approximation
As can be seen from Eqs. (10) and (11), in the leading-order approximation for $\rho\ll 1$, the CD spacetime geometry is characterized by an effective metric that is independent of $\eta$:
@@
Finally, considering once again the approximation $\rho\ll 1$, we find the asymptotic behavior of $\eta$ and $\tau$ as functions of $\rho$:
@@
Analogously to the previous case, the CD spacetime geometry at leading order approximation for $\rho\gg 1$ exhibits an effective metric that is independent of $\eta$ (see Eqs. (14) and (15)):
@@
Interestingly, by defining $4\tilde{\rho}=\alpha^{2}/\rho$, which corresponds to the regime $\tilde{\rho}\ll 1$, the metric reduces to the same form as that obtained in the case $\rho\ll 1$:
@@
Proceeding analogously to the case of $\rho\ll 1$, we find that in the regime $\tilde{\rho}\ll 1$, the asymptotic behavior of $\eta$ and $\tau$ as a function of $\tilde{\rho}$ is given by:
@@ III Field Modes in the Near-Horizon Approximation and Conformal Quantum Mechanics
As discussed in Sec. II.2, around the regimes $\rho\ll 1$ and $\rho\gg 1$, the conformal factor $\Lambda^{2}$ exhibits the following asymptotic behavior:
@@
Thus, in both asymptotic regimes, $\rho\ll 1$ and $\rho\gg 1$, the field equation reduces to a Schrödinger-like eigenvalue problem characteristic of conformal quantum mechanics (CQM), a one-dimensional theory with $\mathfrak{sl}(2,\mathbb{R})$ symmetry (see Ref. [21]).
@@
where $(+)$ corresponds to outgoing modes (away from the horizon) and ()(-) to ingoing modes (towards the horizon). Consequently, the near-horizon field modes with positive frequency take the form of $(\pm)$ CQM modes:
@@ IV.1 Setup: Atom–Field Interaction
The physical system under consideration consists of a quantum scalar field and a cloud of two-level atoms, interacting weakly along the atoms’ trajectories. This setup was originally proposed in Ref. [1], which introduced the phenomenon of Horizon Brightened Acceleration Radiation (HBAR). Here, we apply this framework to the CD spacetime geometry, which distinguishes our study from previous works focusing on black hole geometries [9, 11, 12, 13].
@@
The other component of the system consists of a cloud of atoms, modeled as two-level systems. In this work, we focus on the process in which atoms are injected at random times from the asymptotic past, $\eta\to$-$\infty$, at $\rho\gg 1$, and subsequently follow geodesic trajectories toward the causal horizon $\mathcal{H}^{+}_{0}$ (the analogous injection process from $\eta\to$-$\infty$ at $\rho\ll 1$, directed toward $\mathcal{H}^{+}_{\infty}$, is discussed in Appendix A). Being massive, the atoms eventually reach the asymptotic future, $\eta\to\infty$, terminating at $\rho\ll 1$ within the region $i^{+}$, close to the left corner, as illustrated in Fig. 6(b). Furthermore, while moving along these free-fall trajectories, the atoms interact with the quantum field. Although the field remains in its vacuum state, the freely falling atoms perceive it differently due to their relative acceleration with respect to the field modes. As a consequence, the exchange of virtual scalar photons is disrupted, leading to spontaneous emission and absorption processes in which the atoms interact with scalar field quanta (see Refs. [1, 12]).
The interaction between each atom and the field is modeled by the dipole interaction Hamiltonian in the interaction picture:
@@ IV.2 Emission and absorption probabilities
Given our interest in characterizing the physical nature of these processes, it is more convenient to compute the corresponding rates of emission and absorption, defined as $R_{\rm em,\mathbf{s}}=\mathfrak{r}P_{\rm em,\mathbf{s}}$ and $R_{\rm ab,\mathbf{s}}=\mathfrak{r}P_{\rm ab,\mathbf{s}}$, respectively. Here, $\mathfrak{r}=\Delta N/\Delta\eta$ denotes the atom injection rate, with $\Delta N$ representing the number of atoms injected during the interval $\Delta\eta$ (see Ref. [12] for further details). In the following, we focus on evaluating the emission rate near the causal horizon $\mathcal{H}^{+}_{0}$ (i.e., for $\rho\ll 1$), analyzing the integral characterizing this process as given in Eq. (44) for the outgoing $(+)$ CQM mode:
@@
The behavior of this integral is governed by the interplay between the two oscillatory factors, $\rho\,\rho^{-2i\omega/\kappa}$ and $e^{-i8\nu\kappa\rho^{2}/\tilde{e}}$. The first factor can be decomposed into a logarithmically oscillating term, $\rho^{-2i\omega/\kappa}$—which originates from the conformal quantum mechanics (CQM) structure of the near-horizon region—and a linear prefactor $\rho$, which suppresses the amplitude while remaining finite for $\rho>0$. The scale invariance of $\rho^{-2i\omega/\kappa}$, characterized by the dimensionless conformal parameter $\Theta=\omega/\kappa$, makes it the dominant contribution in the regime $\rho\ll 1$, where the exponential factor $e^{-i8\nu\kappa\rho^{2}/\tilde{e}}$ is approximately unity. As $\rho$ increases away from the horizon, the exponential term induces rapid oscillations. Thus, these oscillations average out and contribute negligibly to the integral. Consequently, the upper limit $\rho_{\circ}$ in Eq. (46) can be effectively extended to infinity:

### v2

# Horizon brightened acceleration radiation entropy in causal diamond geometry: A near-horizon perspective

@@ Abstract
In this article, we extend the horizon brightened acceleration radiation (HBAR) framework, originally introduced by Marlan Scully et al. in Proceedings of the National Academy of Sciences 115, 8131 (2018), to the causal diamond (CD) spacetime. We study a cloud of two-level atoms, injected at random times in the asymptotic past of the CD, freely falling toward its causal horizon and emitting scalar radiation via a weak dipole coupling to a quantum field. In the near-horizon region, an emergent conformal symmetry—captured by conformal quantum mechanics (CQM)—governs the field dynamics and allows analytic control of the emission process. We find that the radiation spectrum is thermal, with temperature $T_{D}=1/(\pi\alpha)$, and that the associated von Neumann entropy flux reproduces the entropy production of the radiation field. These results demonstrate that the causal horizons of the CD spacetime effectively act as a topological thermal reservoir, with thermal properties arising entirely from the global causal structure rather than from underlying microscopic degrees of freedom, highlighting that the validity of the HBAR framework is fundamentally tied to the existence of causal horizons, independent of the presence of a black hole.
@@ I Introduction
In the context of horizon brightened acceleration radiation (HBAR) entropy, as proposed by Marlan Scully et al. in Ref. [1], it has been argued that the nature of the radiation emitted by a cloud of atoms—randomly injected from the asymptotic past and following free-falling (geodesic) trajectories toward the event horizon of a black hole—is a consequence of the relative acceleration between the field modes, defined on the black hole geometry with a well-defined Boulware vacuum state [2], and the locally inertial motion of the atoms. The origin of this radiation lies in the weak dipole coupling between the atoms, modeled as two-level systems, and the quantum field [3]. A characteristic process of this interaction is the excitation of an atom accompanied by the emission of a scalar quantum, which can either reach future timelike infinity—thereby contributing to the observable radiation spectrum—or be reabsorbed by another atom within the cloud before being emitted again.
@@
Consequently, the HBAR framework finds a natural extension in the near-horizon geometry, providing a powerful tool for analyzing more complex spacetime geometries, especially in situations where an exact geodesic analysis is intractable. The conformal symmetry that emerges in this regime allows for analytic control over the emission and absorption spectra of scalar quanta associated with acceleration radiation. This approach has proven fruitful in subsequent analyses of HBAR entropy for a variety of geometries, including Schwarzschild black holes [9], Kerr black holes [11, 12, 13], quantum-corrected black hole geometries [14, 15, 16], along null geodesics [17], braneworld black holes [18], and analyses involving pointlike and finite-size detectors [19].
@@ II.1 Coordinate transformation and metric tensor
The spacetime experienced by such a finite-lifetime observer defines the so-called causal diamond (CD) spacetime. It can be geometrically characterized from the subset of Minkowski spacetime points restricted to the region $D_{R}$, denoted as $x^{\mu}=(t_{d_{R}},x_{d_{R}})$. These coordinates are obtained by applying the conformal transformation introduced in Ref. [20] to the Minkowski coordinates restricted to the right Rindler wedge $R$ [see Fig. 1(b)], which are given by $\tilde{x}^{\mu}=(t_{r},x_{r})$. Accordingly, the conformal map is defined as
@@
where $\tilde{\alpha}=2\alpha/\lambda$. Then, by applying Eqs. (1)–(5), we obtain the coordinate transformation between the Minkowski points restricted to the region $D_{R}$ and the CD spacetime coordinates $(\eta,\rho)$
@@
The line element of the CD spacetime takes the form
@@ II.2.1 Causal diamond horizons
We begin by analyzing the expansion of $(t_{d_{r}},x_{d_{r}})$ around the regime $\rho\ll 1$, which is given by
@@
In addition, since $\sinh(2\eta/\alpha)\approx\pm\cosh(2\eta/\alpha)$ for $\eta\gg 1$ and $\eta\gg 1$, respectively, it follows from Eqs. (10) and (11) that, in the simultaneous limits $\rho\to 0$ and $\eta\to\pm\infty$, one directly obtains the following causal horizons [see Fig. 3(a)]
@@
On the other hand, we now turn to the expansion of $(t_{d_{r}},x_{d_{r}})$ in the regime $\rho\gg 1$, from which we obtain
@@
As in the previous case, applying the limits $\rho\to\infty$ and $\eta\to\pm\infty$ to Eqs. (14) and (15) yields the following causal horizons [see Fig. 3(b)]
@@ II.2.2 Causal diamond corners
The asymptotic regions of the CD spacetime corresponding to the four corners of $D_{R}$ can be identified by analyzing the coordinate transformations between Minkowski and the CD spacetime in the appropriate limits
@@ II.3 Effective surface gravity of the CD spacetime horizons
Likewise, noting that the boundary of the CD spacetime corresponds, in Minkowski spacetime, to the boundary of the region $D_{R}$, given by $\partial D_{R}:=\quantity{(t,x):\lvert t\rvert+\lvert x\rvert=\alpha}$, we observe that it precisely corresponds to the Killing horizon of the vector field $\chi^{(\eta)}$
@@
Thus, the effective surface gravity associated with the conformal Killing vector $\chi^{(\eta)}$ is defined by evaluating the following identity on the boundary of $D_{R}$
@@
where $\hat{\epsilon}=\pm 1$ depending on whether the normal vector to a given segment of the boundary of $D_{R}$ is future-directed $(+)$ or past-directed ()(-) [see Fig. 5(b)].
Finally, by evaluating Eq. (24) explicitly in Minkowski coordinates—using the expression for $\chi^{(\eta)}$ given in Eq. (22) together with the relation $\chi^{(\eta)}\cdot\chi^{(\eta)}=\eta_{ab}\chi^{a}_{(\eta)}\chi^{b}_{(\eta)}$—we obtain the following value for the effective surface gravity and its corresponding Hawking-like temperature
@@
Thus, the obtained temperature exactly matches the temperature $T_{D}$ associated with the thermal description of the Minkowski vacuum as perceived by an observer with finite lifetime $2\alpha$, restricted to the causal diamond $D_{R}$. This result is consistent with previous findings based on the thermal time hypothesis, the open quantum system approach, the thermofield double construction, and related frameworks [23, 24, 25, 26, 27, 28, 29]. In what follows, we adopt this identification of $T_{D}$ in terms of the surface gravity $\kappa$, namely $T_{D}=\kappa/(2\pi)=(\pi\alpha)^{-1}$.
@@ II.4.1 Geodesic trajectories
From the preceding geometric analysis, we conclude that, when viewed from Minkowski spacetime, the causal diamond $D_{R}$ exhibits a structure analogous to that of a Penrose-Carter diagram. This correspondence follows from the fact that the CD spacetime metric, given in Eq. (8), is conformally related to the Minkowski metric through the coordinate transformation $\rho=e^{2\xi/\alpha}$.
@@
Figure 6: (a) timelike geodesics. (b) a timelike geodesic in the extended representation of $i^{-}$ and $i^{+}$. (c) Null geodesics.
@@ II.4.2 Geodesics in the near-horizon approximation
As can be seen from Eqs. (10) and (11), in the leading-order approximation for $\rho\ll 1$, the CD spacetime geometry is characterized by an effective metric that is independent of $\eta$
@@
Finally, considering once again the approximation $\rho\ll 1$, we find the asymptotic behavior of $\eta$ and $\tau$ as functions of $\rho$
@@
Analogously to the previous case, the CD spacetime geometry at leading order approximation for $\rho\gg 1$ exhibits an effective metric that is independent of $\eta$ [see Eqs. (14) and (15)]
@@
Interestingly, by defining $4\tilde{\rho}=\alpha^{2}/\rho$, which corresponds to the regime $\tilde{\rho}\ll 1$, the metric reduces to the same form as that obtained in the case $\rho\ll 1$
@@
Proceeding analogously to the case of $\rho\ll 1$, we find that in the regime $\tilde{\rho}\ll 1$, the asymptotic behavior of $\eta$ and $\tau$ as a function of $\tilde{\rho}$ is given by
@@ III Field Modes in the Near-Horizon Approximation and Conformal Quantum Mechanics
As discussed in Sec. II.2, around the regimes $\rho\ll 1$ and $\rho\gg 1$, the conformal factor $\Lambda^{2}$ exhibits the following asymptotic behavior
@@
Thus, in both asymptotic regimes, $\rho\ll 1$ and $\rho\gg 1$, the field equation reduces to a Schrödinger-like eigenvalue problem characteristic of conformal quantum mechanics (CQM), a one-dimensional theory with $\mathfrak{sl}(2,\mathbb{R})$ symmetry (see Ref. [30, 31, 32]).
@@
where $(+)$ corresponds to outgoing modes (away from the horizon) and ()(-) to ingoing modes (toward the horizon). Consequently, the near-horizon field modes with positive frequency take the form of $(\pm)$ CQM modes
@@ IV.1 Setup: Atom–field interaction
The physical system under consideration consists of a quantum scalar field and a cloud of two-level atoms, interacting weakly along the atoms’ trajectories. This setup was originally proposed in Ref. [1], which introduced the phenomenon of HBAR. Here, we apply this framework to the CD spacetime geometry, which distinguishes our study from previous works focusing on black hole geometries [9, 11, 12, 13].
@@
The other component of the system consists of a cloud of atoms, modeled as two-level systems. In this work, we focus on the process in which atoms are injected at random times from the asymptotic past, $\eta\to$-$\infty$, at $\rho\gg 1$, and subsequently follow geodesic trajectories toward the causal horizon $\mathcal{H}^{+}_{0}$ (the analogous injection process from $\eta\to$-$\infty$ at $\rho\ll 1$, directed toward $\mathcal{H}^{+}_{\infty}$, is discussed in Appendix A). Being massive, the atoms eventually reach the asymptotic future, $\eta\to\infty$, terminating at $\rho\ll 1$ within the region $i^{+}$, close to the left corner, as illustrated in Fig. 6(b). Furthermore, while moving along these free-fall trajectories, the atoms interact with the quantum field. Although the field remains in its vacuum state, the freely falling atoms perceive it differently due to their relative acceleration with respect to the field modes. As a consequence, the exchange of virtual particles is disrupted, leading to spontaneous emission and absorption processes in which the atoms interact with scalar field quanta (see Refs. [1, 12]).
The interaction between each atom and the field is modeled by the dipole interaction Hamiltonian in the interaction picture
@@ IV.2 Emission and absorption probabilities
Given our interest in characterizing the physical nature of these processes, it is more convenient to compute the corresponding rates of emission and absorption, defined as $R_{\rm em,\mathbf{s}}=\mathfrak{r}P_{\rm em,\mathbf{s}}$ and $R_{\rm ab,\mathbf{s}}=\mathfrak{r}P_{\rm ab,\mathbf{s}}$, respectively. Here, $\mathfrak{r}=\Delta N/\Delta\eta$ denotes the atom injection rate, with $\Delta N$ representing the number of atoms injected during the interval $\Delta\eta$ (see Ref. [12] for further details). In the following, we focus on evaluating the emission rate near the causal horizon $\mathcal{H}^{+}_{0}$ (i.e., for $\rho\ll 1$), analyzing the integral characterizing this process as given in Eq. (44) for the outgoing $(+)$ CQM mode
@@
The behavior of this integral is governed by the interplay between the two oscillatory factors, $\rho\,\rho^{-2i\omega/\kappa}$ and $e^{-i8\nu\kappa\rho^{2}/\tilde{e}}$. The first factor can be decomposed into a logarithmically oscillating term, $\rho^{-2i\omega/\kappa}$—which originates from the conformal quantum mechanics (CQM) structure of the near-horizon region—and a linear prefactor $\rho$, which suppresses the amplitude while remaining finite for $\rho>0$. The scale invariance of $\rho^{-2i\omega/\kappa}$, characterized by the dimensionless conformal parameter $\Theta=\omega/\kappa$, makes it the dominant contribution in the regime $\rho\ll 1$, where the exponential factor $e^{-i8\nu\kappa\rho^{2}/\tilde{e}}$ is approximately unity. As $\rho$ increases away from the horizon, the exponential term induces rapid oscillations. Thus, these oscillations average out and contribute negligibly to the integral. Consequently, the upper limit $\rho_{\circ}$ in Eq. (46) can be effectively extended to infinity